OpenStudy (anonymous):

Solve for x: -3|x+7| = -12

jagr2713 (jagr2713):

x=−3,−11

jagr2713 (jagr2713):

you want explanation

OpenStudy (anonymous):

Thanks... so much man, please give me an explanation... I don't truly understand this at all...

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

yeah I can get it for you

jagr2713 (jagr2713):

ok is my answer one of them

OpenStudy (anonymous):

A x = 5 over 3, x = −19 over 3 B x = −3, x = −11 C x = −3, x = 11 D No solution

OpenStudy (anonymous):

Yes yours is one of the answers

jagr2713 (jagr2713):

Divide each term in the equation by 3. Cancel the common factor of 3 in −3|x+7|\3. Reduce the expression −3|x+7|\3 by removing a factor of 3 from the numerator and denominator. Cancel the common factor of 3 in −12/3. Reduce the expression −12/3 by removing a factor of 3 from the numerator and denominator. and you get −|x+7|=−4

jagr2713 (jagr2713):

then

jagr2713 (jagr2713):

Multiply each term in the equation by −1. −|x+7|⋅−1=−4⋅−1 Multiply −|x+7| by −1 to get |x+7|. |x+7|=−4⋅−1 Multiply −4 by −1 to get 4. |x+7|=4

jagr2713 (jagr2713):

|x+7|=4 Remove the absolute value term. This creates a ± on the right-hand side of the equation because |x|=±x. x+7=±(4) Set up the + portion of the ± solution. x+7=4 Move all terms not containing x to the right-hand side of the equation.

jagr2713 (jagr2713):

Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides. x=−7+4 Add 4 to −7 to get −3. x=−3

jagr2713 (jagr2713):

so we got -3

jagr2713 (jagr2713):

Set up the − portion of the ± solution. x+7=−(4) Multiply −1 by the 4 inside the parentheses. x+7=−4 Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides. x=−7−4 Subtract 4 from −7 to get −11. x=−11 The solution to the equation includes both the positive and negative portions of the solution. so we get the answer -3,-11

jagr2713 (jagr2713):

did that help

OpenStudy (anonymous):

Now.... I get what your saying... It did help due to the fact that my teacher didn't cover any of this during my regular school year thus prompting me to take an online course to make up for lost time...

OpenStudy (texaschic101):

I do it a little bit differently, but the answer comes out the same. However, if you understand it the way jagr explained, I do not have to show you the way I do it.

OpenStudy (anonymous):

But thank you for the help with that one....

jagr2713 (jagr2713):

yw